OpenStudy (anonymous):

How fast (metric units) is a ball going after 3.0 seconds if you drop it off a very high building? It would be 9.8 m/s^2 since gravity is a constant speed, right?

OpenStudy (anonymous):

v=v0-gt v0 = 0 v = -gt v = -9.8 * 3 = -29.4 m/s (minus because i assume positive y axis is up)

OpenStudy (anonymous):

acceleration 9.8 m/s^2 means that every second you get 9.8 in velocity

OpenStudy (anonymous):

Wait... so what is the answer to this problem? A kid drops a rock in a dark well. After 3.0 seconds, he hears a splash. How deep in meters is the well? Would that one be 29.4 also?

OpenStudy (anonymous):

no y = wellHeight -gt^2 / 2 0 = wellHeight - 9.8 * (3)^2 / 2 wellHeight = 44.1

OpenStudy (anonymous):

44.1 meters

OpenStudy (calculusfunctions):

Do know the kinematic equations or Galileo's equation for a freely falling body?

OpenStudy (anonymous):

No

OpenStudy (calculusfunctions):

You've never taken a physics course?

OpenStudy (calculusfunctions):

Do you know calculus?

OpenStudy (anonymous):

I'm taking a physics course right now.

OpenStudy (anonymous):

I'm in precal

OpenStudy (calculusfunctions):

Have you learned derivatives?

OpenStudy (anonymous):

No

OpenStudy (calculusfunctions):

Sorry, I just returned. OK so if you haven't learned derivatives, then try the following and tell me if you get the answer. In fact post your solution please so if you've made any careless errors, I can check. final speed = initial speed + acceleration x time I'll be anticipating your response.

OpenStudy (anonymous):

29.4 m/sec = 0 + 9.8 m/sec^2 * 3 sec

OpenStudy (calculusfunctions):

final speed = initial speed + acceleration x time final speed = 0 m/s + (-9.8 m/s²)(3.0 s) final speed = -29.4 m/s ∴ the ball is falling at a rate of 29.4 m/s. Note: In the therefore statement the negative sign is omitted because I said that ... the ball is falling ... "Falling" implies descent which implies a negative direction. However, when substituting into the equation, it is necessary to replace the acceleration with -9.8 m/s², because gravity is a force that acts downward. Do you understand now?

OpenStudy (anonymous):

Yes

OpenStudy (calculusfunctions):

Are my answers not good enough to deserve a medal?

OpenStudy (anonymous):

Haha fine...